# An Event Structure-aware Generative Model for Biomedical Event Extraction

### v3

@@ IV-C Baseline Models
IV-D Experimental Results
Table IV shows the performance results based on the MLEE, GE11, and PHEE datasets. “Gen” and “Cls” denote Generation-based models and Classification-based models, respectively. The letter “l” in the column PLM is used to denote the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, the GenBEE model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we reckon that the simpler event structures with less structural information in the PHEE datasets are the main reason that affects the performance of the GenBEE model.
Table IV also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in the experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of +3.4%, +3.0%, and +2.2% on Trg-C F1-score, and an improvement of +3.3%, +3.9% and +2.4% on Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.
@@ IV-D Experimental Results
V Analysis
V-A Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different modules on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following two modules: Event Prompt Construction and Structural Prefix Learning. Table V reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table V, we observe that event prompts and structural prefixes significantly contribute to the performance of our proposed model, regardless of the dataset. Removing the Event Prompt Construction module leads to a decrease of -0.6% and -1.0% on the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Moreover, the removal of the Structural Prefix Learning module from GenBEE reduces the performance by -1.5% and -1.6% on the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this module in the model’s overall performance. Finally, removing both Event Prompt Construction and Structural Prefix Learning modules will further degrade the model’s performance. Overall, each module of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
TABLE V: Ablation study on the proposed GenBEE model based on different model configurations. We report the results in F1-scores (%).
@@ V-A Ablation Study
V-B Performance Comparison between LLMs API and Fine-tuned PLMs
Given the immense potential of using large language models (LLMs) API with In-Context Learning (ICL) [21] across various NLP tasks under data efficient scenarios [38], we conduct experiments to compare the few-shot performance of the proposed GenBEE model with LLMs API. Specifically, we consider two widely used LLMs: GPT-4 [39] and Llama3-70b [40] for comparison. We access these LLMs through APIs from their official providers. As a part of the prompt, we provide LLMs with the same type of instructions, EE templates that we use in GenBEE, and a few ICL examples (positive ones). It is worth noting that the number of ICL examples is restricted by the maximum context length supported by LLMs.
Table VI presents the performance results for LLMs and our proposed GenBEE model, categorized by shot numbers: 4-shot, 8-shot, 16-shot, and 32-shot. In the 4-shot learning scenario, the GPT-4 API achieves the highest F1 scores across all datasets, followed by the Llama3-70b API, with GenBEE showing the lowest performance. Similar trends are observed in the 8-shot and 16-shot learning scenarios, where all models exhibit performance improvements with increased learning data (shot numbers). However, for LLMs, adding more ICL examples beyond 16-shot does not necessarily result in further performance gains, as improvements tend to plateau or even slightly decline at higher shot levels. Notably, at 32-shot learning, GenBEE achieves the highest F1-scores across all datasets, particularly on the PHEE dataset with F1-scores of 43.7% and 38.5% for Trigger and Argument Classification, respectively. We infer that this is likely due to PHEE’s limited number of event types (only two), which reduces the need for more number of unseen labels during few-shot training and allows GenBEE to leverage the provided examples more effectively. In comparison, datasets such as MLEE and GE11 which contain more diverse event types present a greater challenge for the GenBEE model. Generally, the fine-tuned GenBEE with few-shot learning is able to outperform LLMs (using In-Context Learning) when the number of labeled data increases.
@@ V-B Performance Comparison between LLMs API and Fine-tuned PLMs
TABLE VI: Experimental results using LLMs API and GenBEE with few-shot learning based on the MLEE, GE11 and PHEE datasets. All reported results are in F1-scores (%).
Models MLEE GE11 PHEE
Trg-C Arg-C Trg-C Arg-C Trg-C Arg-C
4-shot
GPT-4 API 34.0 27.5 23.0 21.9 39.5 33.8
Llama3-70b API 31.8 26.7 19.3 20.5 36.5 30.5
GenBEE 11.1 8.4 5.3 5.5 21.0 14.5
8-shot
GPT-4 API 35.7 29.0 24.1 22.1 40.4 34.5
Llama3-70b API 33.0 28.5 20.3 20.7 37.4 31.8
GenBEE 18.0 14.2 9.0 7.9 29.8 24.6
16-shot
GPT-4 API 34.4 30.2 24.7 22.8 41.7 35.2
Llama3-70b API 33.2 30.0 20.6 22.0 37.8 32.0
GenBEE 27.6 22.8 18.4 14.3 38.6 33.7
32-shot
GPT-4 API 34.2 29.5 23.9 23.2 41.2 35.4
Llama3-70b API 32.9 29.8 20.5 21.8 36.9 31.5
GenBEE 35.7 31.4 25.2 24.3 43.7 38.5

### v5

@@ 4.1 Experimental Setup
Metrics. For evaluation metrics, we follow the previous work [34, 25, 8] to adopt the following two metrics to evaluate the performance: (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
Implementation Details. For a fair comparison with recent works, we leverage the BART-large [17] as the PLM in our proposed GenBEE model. For encoding structural prompts, we use a pre-trained BioBERT [24] model. We train GenBEE on an NVIDIA A100 40G GPU. The learning rate of BART-large is set to $10^{-5}$ and the learning rate of BioBERT is set to $10^{-6}$. Moreover, we train GenBEE for 50 epochs on PHEE and 80 epochs on MLEE and GE11. The batch size is set to 16 during training. The prefix length $l$ is set to 40, which is determined experimentally within the interval $l$ = {20, 30, 40, 50, 60}.
Baseline Models. We compare our GenBEE with the following EE models: (1) DyGIE++ [34] is a classification-based model that captures contextual information with the span graph propagation technique. (2) OneIE [22] is a classification-based model that utilizes global feature-aware graphs to capture cross-subtask and cross-instance interactions. (3) AMR-IE [38] is a classification-based model that captures syntactic characteristics in contexts using the AMR graph. (4) EEQA [4] is a classification-based model that formulates EE as a question-answering task. (5) TagPrime [9] is a state-of-the-art classification-based EE model that iteratively extracts event information based on each event type. (6) DEGREE [8] is a state-of-the-art generation-based EE model that integrates prompts for conditional generation. As the vanilla DEGREE model was designed mainly for general domain event extraction, we have re-implemented it using the prompts provided by [12] for the MLEE, GE11, and PHEE datasets. To ensure a fair comparison across models, we adopt the official codes of the above baselines and train them with the same data.
4.2 Experimental Results
Table 4 shows the performance results based on the MLEE, GE11, and PHEE datasets. “Gen" and “Cls" denote Generation-based models and Classification-based models, respectively. The letter “l” in the column PLM is used to denote the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, the GenBEE model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we observe that on the PHEE dataset, which only contains few nested and overlapping event structures, GenBEE exhibits a marginal reduction in performance.
Table 4 also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in the experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of +3.4%, +3.0%, and +2.2% on Trg-C F1-score, and an improvement of +3.3%, +3.9% and +2.4% on Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.
@@ 4.2 Experimental Results
Table 5: Ablation study on the proposed GenBEE model based on different model configurations. We report the results in F1-scores (%).
@@
4.3 Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different modules on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following two modules: Event Prompt Construction and Structural Prefix Learning. Table 5 reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table 5, we observe that event prompts and structural prefixes significantly contribute to the performance of our proposed model, regardless of the dataset. Removing the Event Prompt Construction module leads to a decrease of -0.6% and -1.0% on the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Moreover, the removal of the Structural Prefix Learning module from GenBEE reduces the performance by -1.5% and -1.6% on the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this module in the model’s overall performance. Finally, removing both Event Prompt Construction and Structural Prefix Learning modules will further degrade the model’s performance. Overall, each module of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
@@ 4.3 Ablation Study
Figure 4: Experimental results using LLMs and GenBEE with few-shot learning based on the MLEE and GE11 datasets. All reported results are in Arg-C F1-scores (%).
4.4 Few-shot Learning Performance Comparison with LLMs
Given the immense potential of using large language models (LLMs) API with In-Context Learning (ICL) [3] across various NLP tasks under data efficient scenarios [36], we conduct experiments to compare the few-shot performance of the proposed GenBEE model with LLMs. Specifically, we consider two widely used LLMs: GPT-4 [1] and Llama3-70b [32] for comparison. We access these LLMs through APIs from their official providers. As a part of the prompt, we provide LLMs with the same type of instructions, EE templates that we use in GenBEE, and a few ICL examples (positive ones). It is worth noting that the number of ICL examples is restricted by the maximum context length supported by LLMs.
Figure 4 shows the few-shot learning performance results for LLMs and our proposed GenBEE model across the MLEE and GE11 benchmarks. The results indicate that while GPT-4 and Llama3-70b demonstrate significant advantages in extremely data-scarce scenarios, such as 4-shot and 8-shot settings, GenBEE begins to close the performance gap as the number of labeled examples increases, and eventually outperforms the LLMs. For example, in the 32-shot and 64-shot settings, the GenBEE model surpasses both LLMs. Generally, while LLMs with In-Context Learning are well-suited to extremely data-scarce scenarios, the fine-tuned GenBEE is able to outperform LLMs as the amount of labeled data increases.
@@ 5 Conclusion
Acknowledgment
This work was enabled in part by funding from the National Science Foundation awards: 2201428, 2232862, 2004014, 2138259, and 2406251. We appreciate valuable suggestions from the anonymous reviewers.